# In-context learning of evolving data streams with tabular foundational models

### v1

@@
Abstract
State-of-the-art data stream mining in supervised classification has traditionally relied on ensembles of incremental decision trees. However, the emergence of large tabular models, i.e., transformers designed for structured numerical data, marks a significant paradigm shift. These models move beyond traditional weight updates, instead employing in-context learning through prompt tuning. By using on-the-fly sketches to summarize unbounded streaming data, one can feed this information into a pre-trained model for efficient processing. This work bridges advancements from both areas, highlighting how transformers’ implicit meta-learning abilities, pre-training on drifting natural data, and reliance on context optimization directly address the core challenges of adaptive learning in dynamic environments. Exploring real-time model adaptation, this research demonstrates that TabPFN, coupled with a simple sliding memory strategy, consistently outperforms ensembles of Hoeffding trees across all non-stationary benchmarks. Several promising research directions are outlined in the paper. The authors urge the community to explore these ideas, offering valuable opportunities to advance in-context stream learning.
Keywords Data stream  $\cdot$ concept drift  $\cdot$ transformer  $\cdot$ in-context learning
1 Introduction
@@
Data stream mining is an area in machine learning where the inference and training of the algorithm are performed in real-time, dealing with large volumes of ever-evolving tabular data. Differently to batch learning where all training data necessary to induce a model is available, data streams incrementally arrive at any time. Thus, making a model easily outdated due to the occurrence of concept drifts, i.e., distribution changes over time [1]. To circumvent these challenges, many stream mining algorithms have been proposed with different requirements being imposed: (1) be ready to predict and update the model at any point, and in limited time; (2) process an example at a time, inspecting only once with a limited sized model; (3) be able to adapt to change.
@@
Decision trees. In fulfilling these requirements, the current state-of-art for supervised learning in tabular data streams has long been incremental decision trees (IDTs), whose success can be attributed to two key factors: approximation-based splitting techniques [2] and effective adaptation strategies [3]. Approximation-based splitting involves updating statistical summaries of entropy-based metrics, e.g. information gain, and determining whether the observed utility of a split is statistically close to its true utility when the data distribution is unknown, e.g. via the Hoeffding bound [4]. However, as tree grows from the root node, the descendant nodes subsequently get fixed to covering particular sub-spaces of the space covered by their parent node. To address this, adaptation strategies allow to trigger the forgetting of old knowledge, pruning the affected parent nodes, instead of a complete replacement of the base learner [5]. Alternatively, multiple trees can be combined into a high diversity ensemble, which naturally develops decision boundaries that locally adapt to changes in the data distribution [6, 7]. Upon a concept drift, the diversity among trees can be exploited as some local minima are less affected than others. Affected components can be dynamically deleted, added, and combined with a voting scheme, weighting different weak learners based on their past performance.
Lack of autonomy. Regardless of how sophisticated these strategies are, these streaming algorithms still rely on human expertise to select the most suitable algorithm, or determine the optimal configuration for a given algorithm. Similarly to the way drift detectors identify when the specific trained model is no longer performing, the model class itself may no longer be applicable. The performance curves of different streaming algorithms, inductive biases, and configurations cross as the stream evolves. In these moments, the learning framework should autonomously select both the right algorithm [8] and corresponding hyperparameter settings [9].
Large Language Models (LLMs). Conversely, the disruptive success and knowledge-modeling capabilities of LLMs, e.g. GPT-3 [10] and LLaMA [11] brings the opportunity to tap into prior experience, to the point that most of the required learning has already been done beforehand. While the aforementioned streaming methods follow the traditional two-stage methodology, i.e., first optimizing the model with several tuning iterations over a sliding window to then select the best algorithm or hyperparameter configuration for the next window, LLMs allow to instantaneously deploy a model, bypassing the computational and temporal costs of this traditional process. By pretraining on extremely large corpora of textual datasets via autoregressive next-token prediction, LLMs develop common sense inductive biases from the infinite amounts of prior learning experiences. Thus, driving emergent abilities that improve with scale, such as zero- and few-shot in-context learning (ICL), which allows models to perform new tasks at inference time by conditioning on a prompt of input-output examples, without any parameter updates [10, 12].
Tabular prompts. Consequently, this ICL ability pushed a new research paradigm, focused on adapting pre-trained models through prompt tuning rather than adjusting model weights, by discrete search with natural language, or optimizing the model’s embedding space directly. Within this paradigm and the emerging possibilities of multimodal applications, various efforts were made for flexible prompts of non-textual information, such as structured tabular data tasks with heterogeneous feature spaces and mixed-type columns [13]. As LLMs’ capabilities on table understanding were validated, new efforts went towards directly applying them for supervised learning of tabular data, i.e. predicting an unseen sample by continuing its textual description on the target column [14, 15]. Despite the potential of these advancements to perform semantically-driven drift detection and model adaptation, several limitations are expected to persist in the near term hindering their application for data stream mining: limited context windows, quadratic complexity costs, struggle with autoregressively understanding continuous variables [16, 17, 18], and sensitivity to unexpected characters [19, 20].
Large Tabular Models (LTMs). To alleviate all these limitations, one can shift to designing a dedicated architecture for numeric data from scratch, completely pre-trained on a wide range of tabular datasets with different data distributions, allowing it to learn relevant and general meta-features. These are referred to as LTMs [21]. While it has been known for a while that transformers can handle variable length input sequences in order to perform pretraining and cross-table knowledge transfer [22], it has been recently demonstrated that finetuning from a pretrained tabular transformer [23, 24, 25] is superior to training tabular transformers from scratch [26]. Following these findings, various LTMs pretrained on synthetic tabular data have been proposed for classification [27, 28, 29, 30] and time series forecasting [31, 32]. For instance, the original TabPFN [33] is pre-trained with a 12-layer Transformer for 18000 batches of 512 synthetically generated datasets generated using randomly initialized neural networks to impose the diverse interrelations, e.g. causal relationships and feature importance, that exist among the features of realistic tabular datasets [34]. While this training step is moderately expensive, requiring a total of 20 hours on 8 GPUs (Nvidia RTX 2080 Ti), it is done offline, in advance, and only once, as part of the algorithm development. Then, in deployment TabPFN performs instant classification without fine-tuning, adapting to unseen datasets in a single forward pass at inference time, by using various training examples as context, analogously to how LLMs use the preceding tokens [10]. However, in contrast to LLMs, LTMs like TabPFN allow scaling to much larger datasets, saving tuning resources during deployment, while providing consistent behavior supported by traditional statistical learning.
This work. Building on these advancements, this work proposes a new paradigm, using on-the-fly techniques to summarize unbounded data streams before feeding them to LTMs. While typical streaming methods rely on adapting the model parameters to deal with non-stationary data distributions, this new paradigm uses input data as learnable parameters without changing model parameters, decoupling the design of the learning algorithm from the foundational model. The main idea is to transform big data into a sketch whose size and running time has little or no dependency on the size of the data stream, as illustrated in Figure 1. The learning algorithm is then applied to this dynamically contextualized sketch, with the main challenge of balancing the trade-off between sketch size and information loss. Thus, shifting from in-weights learning to in-context learning. The contributions are outlined as follows:
  • Section 2 provides an overview of the evolution of state-of-the-art methods for ensembles of IDT and transformers in tabular classification tasks.
  • Section 3 offers insights into the drivers behind ICL ability and emphasizes the importance of locality for generalizing effectively to an evolving data stream. It discusses the management of biases introduced by both the foundational model and in-context examples, especially in the presence of concept drift.
  • Section 4 presents experiments on several benchmark tabular datasets, demonstrating that a naive implementation of a sliding short- and long-term memory, paired with a state-of-the-art LTM, consistently outperforms all ensembles of IDT, with a reasonable processing time per instance.
  • Section 5 highlights avenues for future research, including hypernetwork-based LTMs, on-the-fly context optimization strategies, and unique streaming challenges. The authors, too, are interested in exploring these directions further and invite collaboration.
2 Background
This section describes the different advancements for decisions trees and transformers that make them now cross roads as competing solutions for data stream mining. First, describing the different adaptations needed to make ensembles of IDTs the state-of-the art for tabular data stream mining. Second, the different efforts of making transformers capable on working with tabular data and perform ICL.
2.1 Incremental decision trees
Dynamic structural expansion. To incrementally construct a IDT, using contained memory and time per sample, the fundamental algorithm design component is approximation-based splitting [2]. As new instances arrive, they traverse the tree from the root to a leaf node, updating statistics at each node that guarantee for split decisions almost identical to those that conventional batch learning methods would make, despite lookahead and stability issues. Based on these sufficient statistics, IDTs continuously update the heuristic values for each attribute, with a distribution of potential split numeric values being approximate by a single Gaussian [35] and successfully executed according to a statistical Bound of the heuristic difference between attributes [36]. However, while effective for incremental adaptation, this design component does not control for unnecessary tree growth. Thus, posing the risk of: (1) excessive memory consumption, due to multiple redundant splits on all features and increasing running statistics to keep updated; and (2) loss of plasticity, due to interference between concepts and descendant nodes getting subsequently fixed to the space covered by their parent node. To address this, various pre-pruning and pos-pruning techniques have been proposed. Pre-pruning strategies include splitting rules enhancements, adaptive tie breaking, adaptive grace periods, and activity-based expansion modes [37]. Pos-pruning includes various methods that estimate in each decision node whether the corresponding rooted subtree is consistent with the current data and worth pruning, e.g. checking if the current split outperforms the null attribute [38] or directly monitoring the error rate in nodes, e.g. using the Page–Hinckley [39] which is designed to detect abrupt changes in the average of a Gaussian signal [40], or the ADWIN change detector [5]. To reconstruct the tree, one can simply require another attribute to become the top split attribute [38], restructure the otherwise pruned subtree [41], or train an alternative subtree as soon as an inconsistent node is detected, which only eventually replaces the original subtree when its accuracy is superior, and after a user-defined number of examples [5].
Diversity. While single IDTs need necessarily to be pruned, combining multiple trees into a high diversity ensemble allows to naturally develop decision boundaries that locally adapt to changes in the data distribution [42, 7]. Moreover, ensemble strategies can be combined with change detection methods to dynamically select, delete, add, and combine IDTs optimized for different regions of the solution space. For instance, Adaptive Random Forest (ARF) [43] combines resampling with drift recovery strategies. It uses both online bagging and boosting to train classifiers iteratively, increasing the weight on misclassified instances [44]. If a drift is detected, the background tree becomes the new active tree, and the old one is forgotten. Streaming Random Patches (SRP) [6] combines random subspaces and resampling to leverage diversity among base incremental learners, applying a global subspace strategy that increases diversity across weak learners. Streaming Gradient Boosted Trees (SGBT) [45] resets only some parts of the booster using an internal tree replacement mechanism, instead of externally monitoring each item in the boosting ensemble.
Concept history. However, such ensembles constantly train all active base classifiers, progressively discarding some of them when a drift is detected. Base learners evolve and the previously learned concepts are forgotten before they reoccur, if that benefits the current concept being learned. Consequently, these models might need to learn old concepts from scratch, which results in a waste of computational resources, longer training times, and significant prediction errors while models are not up-to-date with the latest state of the data stream. To circumvent this issue, meta-learning [46] and model-based clustering [47] allows to more explicitly model history mechanisms to retain a pool of previous active and inactive concepts. Consequently, maintaining this concept history creates the need of identifying what concept is present at each time. For this purpose, one can assume that two classifiers describe the same concept when predicting similarly during a time window [46], while modeling time dependencies according to the concept drift properties [48]. Alternatively, one can recognize similar concepts based on cohesion and separability in the input space, e.g. euclidean distances between clusters representing different concept clusters [47]. In these, micro-clusters or latent features are used to make a synopsis of the incoming instances and reduce the computational cost of finding similarities among conceptual vectors [49]. Then, methods differ in how they manage this concept history.
@@ 2.1 Incremental decision trees
Algorithm selection. Despite these sophisticated approximation-based splitting, adaptation and concept history strategies, these streaming algorithms still rely on human expertise to select the most suitable algorithm or determine the optimal configuration for a given algorithm. Indeed, combining all these building blocks in a single solution normally yields more inductive biases to assume and hyperparameters to tune. Consequently, as new concepts emerge, the performance curves of different algorithm settings cross. To address this, one can simply build heterogeneous ensembles with different algorithms, whose weights are dynamically selected based on recent performance [8]. Alternatively, one can calculate metafeatures over sliding windows to build a meta-model that predicts the best algorithm [50] or the best drift detector [51] for the next interval of unseen data. Moreover, one can combine both approaches, with the meta-learner choosing an ensemble over the current data [52]. Regarding the hyperparameter settings, one can restart classifier hyperparameter tuning after concept drift [9], meta-optimize drift detectors and classifier hyperparameters on-the-fly [53, 54], or even select the most suitable uncertainty threshold for active learning strategies [55].
2.2 Tabular foundational models
Multi-modal LLMs. Within the discovery of ICL abilities [10], and the subsequent emerging possibilities of multimodal applications, various efforts were made for flexible prompts of tabular information [13]. Initial research efforts focused on complementing multi-modal solutions with table understanding, such as Text2SQL, question answering, fact verification, and natural language inference [56, 57, 58, 59]. For instance, by extending BERT with a masked language model objective and additional embeddings designed to capture tabular structured data [60]. While in simpler architectures, an LLM takes in a query and serialized table as an input prompt, in more involved architectures, the system might be also connected to external spreadsheets, databases or programs, allowing for their manipulation, management, and integration [61, 62, 63].
LLMs for tabular understanding. In order for a foundational model to efficiently perform these tabular tasks it is important to compress tables, not only because large tables might not fit in the context length, but also due to the slow processing of long sentences caused by the quadratic complexity of self-attention [64]. Furthermore, compressing allows to manage the impact of tabular data challenges such as: noisy information in messy data and adversarial inputs [65, 66]; random ordering of rows and table transpositions [67]; arbitrary, sequential and shuffled column names [67]; semi-structured content with rows as strings of key-value pairs and merged contiguous columns [67]; and feature heterogeneity in understanding large tables. Motivated by these challenges, predefined constraints [66] and other methods [60, 68] can be used to truncate the input based on a maximum sequence length. Moreover, the ability to search and retrieve information relevant to the tabular task can also be crucial to automate prompt compression. Instead of manually curating or randomly selecting k examples [62], one can find k examples based on semantic similarity with a query, using a predefined distance metric of the question and example embedding, e.g. the Euclidean or negative cosine similarity [69]. Domain-specific information, such as table and column names can be masked from this query.
LLMs for tabular prediction. Leveraging on these LLMs’ validated capabilities on table understanding, new efforts went towards directly applying them for supervised learning of tabular data. While the same serialization and prompting strategies apply, tabular data prediction requires more sophisticated computation with numerical data, e.g. estimating threshold and densities. To address this, initial methods focused on training on large tabular metadatasets with masked reconstruction loss [56, 57], using text templates to convert tabular data into instruction-oriented formats [70], fine-tuning LLMs on small collections of linearized tabular data to learn fundamental features before adapting it to a specific prediction [15, 71, 72] and fine-tuning on the specific prediction dataset to specialize its knowledge using Low-Rank Adaptation [14].
LLM limitations. Despite the potential of these advancements, several limitations hinder the application of LLMs for tabular supervised learning. Firstly, the limited context window restricts the number of examples that can be utilized for few-shot learning and the amount of additional semantic information that can be ingested, particularly in datasets that contain excessive numbers of categorical columns. Secondly, the high count of parameters, quadratic complexity cost of processing the table as string ($n$ samples · $m$ features) makes inference computationally expensive. Moreover, this computational cost is aggravated in numerical columns, as a single numerical variable is implicitly modeled as an autoregressive series of categorical variables, e.g. 1.23 might require 3 expensive forward calls to the model 1 → . → 23, plus the field separation token that follows. Thirdly, this tokenization and the multinomial training objective impact pattern formation, making LLMs struggle with autoregressively understanding whole numbers and modeling continuous variables, such as simple uniform and Gaussian distributions [16, 17, 18]. Tokenization of integers leads to a fragmented approach where even basic mathematical operations require memorization rather than algorithmic processing. Fourthly, the performance remains sensitive to the precise details of the natural-language input, such as unexpected characters of poorly represented categorical features, inadequate metadata and confusing column names, with no assurances of consistent behavior [19, 20].
Tabular transformers. To address these limitations, one can shift to designing a dedicated architecture for numeric data from scratch [21]. Indeed, the challenge of building transformer-based models for tabular data has been widely studied, using token-based mechanisms for feature selection and reasoning in heterogeneous tabular data [73]. While in text and vision, data is intrinsically tied to the position of a token or pixel, the order-invariance of tabular data makes it challenging to understand the inherent relationships between features, which is further aggravated by heterogeneous and uninformative features, skewed heavy-tailed feature distributions, and extreme values [74, 75]. To circumvent these issues, one can develop a transformer-based architecture using a multi-head self-attentive neural network to automatically learn high-order feature interactions [76], combine column descriptions and cells as input for feature encoding [22], use row and column attention mechanisms to capture the inter-sample interactions with self-supervised pretraining to deal with label scarcity [25], use additive attention to consider the interaction between each token and the global representation, achieving a linear complexity [24], transform categorical features into contextual embeddings [77, 23], or use sequential hierarchical subnetworks that prioritize the most significant features [78, 23].
LTMs. Despite the potential of these methods for pre-training, all these works solely focused on introducing various attention mechanisms to handle the tabular data structure, by attending relevant columns and performing attention across both rows and columns [25], being mostly developed just by training from scratch on each target task. To evaluate the possible ICL capabilities for tabular data of such models, XTab pretrained three of the aforementioned tabular transformers [23, 24, 25] with a diverse range of tabular datasets, separating the models into data-specific and shared components [26]. Thus, finding that finetuning from a pretrained tabular transformer is superior to training tabular transformers from scratch. Following these findings, various LTMs pretrained on synthetic tabular data have been proposed for classification [33, 28, 29, 30] and time series forecasting [31, 32]. In contrast to LLMs, LTMs allow scaling to larger contexts, saving tuning resources during deployment, while providing consistent behavior supported by traditional statistical learning. For instance, TabPFN [33, 27] naturally needs a single floating point number to express a single feature of decimals with two places, compared to the smallest version of LLAMA taking four tokens, i.e. 16384 floating point numbers or a sparse vector of length 128000 [11].
LTMs vs IDTs. Revisiting the aforementioned building blocks of data stream mining with IDTs — dynamic structural expansion, diversity, concept history, and algorithm selection — it becomes evident that LTMs preclude the need to implement these components. Instead of serving as the central elements of the learning process, these building blocks should be embedded within the context provided to the LTM, complementing its already powerful capabilities. In this way, their role shifts from core components to supportive features, ultimately enhancing the LTM’s effectiveness. Furthermore, their implementation aligns with the motivation behind prompt tuning itself, i.e. to improve tabular prediction and scaling with respect to context size (i.e., samples, features, and categories). For this purpose, several strategies have been proposed in batch settings, including summarization [79, 80], meta-representations [30], retrieval-based localized context [81], ensembling [82], and fine-tuning [83, 82]. Further details on how these overlap with a streaming setting will be provided in Section 5.
3 In-context stream learning

### v2

@@
Abstract.
State-of-the-art data stream mining has long drawn from ensembles of the Very Fast Decision Tree, a seminal algorithm honored with the 2015 KDD Test-of-Time Award. However, the emergence of large tabular models, i.e., transformers designed for structured numerical data, marks a significant paradigm shift. These models move beyond traditional weight updates, instead employing in-context learning through prompt tuning. By using on-the-fly sketches to summarize unbounded streaming data, one can feed this information into a pre-trained model for efficient processing. This work bridges advancements from both areas, highlighting how transformers’ implicit meta-learning abilities, pre-training on drifting natural data, and reliance on context optimization directly address the core challenges of adaptive learning in dynamic environments. Exploring real-time model adaptation, this research demonstrates that TabPFN, coupled with a simple sliding memory strategy, consistently outperforms ensembles of Hoeffding trees, such as Adaptive Random Forest, and Streaming Random Patches, across all non-stationary benchmarks.
Concept drift, data stream mining, foundational model
journalyear: 2026copyright: noneconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3780305isbn: 979-8-4007-2258-5/2026/08ccs: Computing methodologies Online learning settingsccs: Computing methodologies Learning under covariate shift
@@ 1. Introduction
Data stream mining is an area in machine learning where the inference and training of the algorithm are performed in real-time, dealing with large volumes of ever-evolving tabular data (gama2010knowledge). Differently to batch learning where all training data necessary to induce a model is available, data streams incrementally arrive at any time. Thus, making a model easily outdated due to the occurrence of concept drifts, i.e., distribution changes over time (gama2014survey). To circumvent these challenges, many stream mining algorithms have been proposed with different requirements being imposed: (1) be ready to predict and update the model at any point, and in limited time; (2) process an example at a time, inspecting only once with a limited sized model; (3) be able to adapt to change.
@@
Decision trees. In fulfilling these requirements, the current state-of-art for supervised learning in tabular data streams has long been incremental decision trees (IDTs), whose success can be attributed to two key factors: approximation-based splitting techniques (lourenco2025dfdt) and effective adaptation strategies (neves2025online). Approximation-based splitting involves updating statistical summaries of entropy-based metrics, e.g. information gain, and determining whether the observed utility of a split is statistically close to its true utility when the data distribution is unknown, e.g. via the Hoeffding bound (domingos2001catching). However, as tree grows from the root node, the descendant nodes subsequently get fixed to covering particular sub-spaces of the space covered by their parent node. To address this, adaptation strategies allow to trigger the forgetting of old knowledge, pruning the affected parent nodes, instead of a complete replacement of the base learner (bifet2009adaptive). Alternatively, multiple trees can be combined into a high diversity ensemble, which naturally develops decision boundaries that locally adapt to changes in the data distribution (gomes2019streaming; krawczyk2017ensemble). Upon a concept drift, the diversity among trees can be exploited as some local minima are less affected than others. Affected components can be dynamically deleted, added, and combined with a voting scheme, weighting learners based on their past performance. Moreover, meta-learning (anderson2016recurring) and model-based clustering (xu2016mining) allows to more explicitly model history mechanisms to retain a pool of previous active and inactive concepts.
@@
Lack of autonomy. Regardless of how sophisticated these strategies are, these still rely on human expertise to select the most suitable algorithm, or determine the optimal configuration for a given algorithm. Indeed, combining all these building blocks in a single solution normally yields more inductive biases to assume and hyperparameters to tune. Consequently, as new concepts emerge, the performance curves of different algorithm settings cross. To address this, one can simply build heterogeneous ensembles with different algorithms, whose weights are dynamically selected based on recent performance (van2015having). Alternatively, one can calculate metafeatures over sliding windows to build a meta-model that predicts the best algorithm (van2014algorithm), the best drift detector (aguiar2023enhancing), or the most suitable uncertainty threshold for active learning strategies (martins2023meta), in the next interval of unseen data.
Foundational models (FMs). Conversely, the disruptive success and knowledge-modeling capabilities of foundational models (brown2020language) brings the opportunity to tap into prior experience, to the point that most of the required learning has already been done beforehand. While the aforementioned streaming methods follow the traditional two-stage methodology, i.e., first optimizing the model with several tuning iterations over a sliding window to then select the best algorithm or hyperparameter configuration for the next window, FMs allow to instantaneously deploy a model, bypassing the computational and temporal costs of this traditional process. By pretraining on extremely large corpora of datasets, FMs develop common sense inductive biases from the infinite amounts of prior learning experiences. Thus, driving emergent abilities that improve with scale, such as few-shot in-context learning (ICL), which allows models to perform new tasks at inference time by conditioning on a set of input-output examples, without any parameter updates (brown2020language).
Textual tabular prompts. Consequently, this ICL ability pushed a new research paradigm, focused on adapting pre-trained models through prompt tuning rather than adjusting model weights, by discrete search with natural language, or optimizing the model’s embedding space directly. Within this paradigm and the emerging possibilities of multimodal applications, various efforts were made for flexible textual prompts of structured tabular data with heterogeneous feature spaces and mixed-type columns (fang2024large). As language-based FMs’ capabilities on table understanding were validated, new efforts went towards directly applying them for supervised learning of tabular data, i.e. predicting an unseen sample by continuing its textual description on the target column (dinh2022lift; hegselmann2023tabllm). However, despite the potential of these advancements to perform semantically-driven drift detection and model adaptation, several limitations are expected to persist in the near term hindering their application for data stream mining: limited context windows, quadratic complexity costs, struggle with autoregressively understanding continuous variables (thawani2021representing; van2024latable), and sensitivity to unexpected characters (zhao2021calibrate).
Large Tabular Models (LTMs). To alleviate all these limitations, one can shift to designing a dedicated architecture for numeric data from scratch, completely pre-trained on a wide range of tabular datasets with different data distributions, allowing it to learn relevant and general meta-features. These are referred to as LTMs (van2024tabular). While it has been known for a while that transformers can handle variable length input sequences in order to perform pretraining and cross-table knowledge transfer (wang2022transtab), it has been recently demonstrated that finetuning from a pretrained tabular transformer (gorishniy2021revisiting; wu2021fastformer; somepalli2021saint) is superior to training tabular transformers from scratch (zhu2023xtab). Following these findings, various LTMs pretrained on synthetic tabular data have been proposed for classification (hollmann2025accurate; bonet2024hyperfast; ye2023training). For instance, the original TabPFN (hollmann2022tabpfn) is pre-trained with a 12-layer Transformer for 18000 batches of 512 synthetically generated datasets generated using randomly initialized neural networks to impose the diverse interrelations that exist among the features of realistic tabular datasets, e.g. causal relationships and feature importance (muller2021transformers). While this training step is moderately expensive, requiring a total of 20 hours on 8 GPUs (Nvidia RTX 2080 Ti), it is done offline, in advance, and only once, as part of the algorithm development. Then, in deployment TabPFN performs instant classification without fine-tuning, adapting to unseen datasets in a single forward pass at inference time by using various training examples as context, analogously to how LLMs use the preceding tokens (brown2020language). However, in contrast to LLMs, LTMs like TabPFN allow scaling to much larger datasets, saving tuning resources during deployment, while providing consistent behavior supported by traditional statistical learning.
This work. Building on these advancements, this work proposes a new paradigm, using on-the-fly techniques to summarize unbounded data streams before feeding them to LTMs. While typical streaming methods rely on adapting the model parameters to deal with non-stationary data distributions, this new paradigm uses input data as learnable parameters without changing model parameters, decoupling the design of the learning algorithm from the foundational model. The main idea is to transform big data into a sketch whose size and running time has little or no dependency on the size of the data stream, as illustrated in Figure 1. The learning algorithm is then applied to this dynamically contextualized sketch, with the main challenge of balancing the trade-off between sketch size and information loss. Thus, shifting from in-weights learning to in-context learning. The main contributions are outlined as follows:
  • Introduce a novel in-context learning paradigm for tabular data streams, formalizing how the distributional qualities of the pre-training data and a mesa-optimization architectural bias address the core challenges of adaptive learning.
  • Extend the popular TabPFN transformer (hollmann2022tabpfn) with a new inference-time sketching mechanism, without altering its pretrained weights. Inspired by a bayesian and frequentistic perspective, we introduce a bias-variance decomposition of prediction error, proving variance reduction through more in-context examples, and addressing drifting context-induced biases via a dual-memory FIFO mechanism that maintains inter-class balance and captures intra-class variability through long- and short-term buffers.
2. In-context stream learning